OpenStudy (anonymous):

Solve: a + 8 - 2 ( a - 12) > 0

OpenStudy (anonymous):

u gotta get rid of parenthesis first so do 2*a and 2* 12

OpenStudy (studygurl14):

a + 8 - 2(a - 12) > 0 a + 8 - 2a + 24 >0 -a + 32 > 0 -a > -32

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Can u help me on a couple more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve: 3(z+1) + 4 ( z - 3 )=5

OpenStudy (anonymous):

3z + 3 + 4z - 12 = 5 1z + 15 = 5 I think..

OpenStudy (anonymous):

Some of my answers are 4/7, 3/7, 1, 2

OpenStudy (studygurl14):

3(z + 1) + 4(z - 3) = 5 3z + 3 + 4z - 12 = 5 7z - 9 = 5 7z = 14 z = 2